OpenStudy (anonymous):

if an atom of a element has a mass number of 31 and 16 neutrons in its nucleus, what is the atomic number of the element?

OpenStudy (anonymous):

Mass no. = No. of Protons + No. of neutrons find no. of protons which is equal to the atomic no.

OpenStudy (anonymous):

so 31 + 16 is my answer?

OpenStudy (anonymous):

31=Z+16 Z=31-16

OpenStudy (anonymous):

yup ryt...the element is phosphorous!